Depreciation for Partial Periods Bean Delivery Company purchased a new delivery truck for $35,400 on April 1, 2016. The truck is expected to have a service life of 5 years or 157,200 miles and a residual value of $4,560. The truck was driven 9,000 miles in 2016 and 10,300 miles in 2017. Bean computes depreciation to the nearest whole month.

Compute depreciation expense for 2016 and 2017 using the
For interim computations, carry amounts out to two decimal places. Round your final answers to the nearest dollar.

  1. Straight-line method
    2016 $
    2017 $
  2. Sum-of-the-years'-digits method
    2016 $
    2017 $
  3. Double-declining-balance method
    2016 $
    2017 $
  4. Activity method
    2016 $
    2017 $
    2. For each method, what is the book value of the machine at the end of 2016? At the end of 2017?
    (Round your answers to the nearest dollar.)
  5. Straight-line method
    2016 $
    2017 $
  6. Sum-of-the-years'-digits method
    2016 $
    2017 $
  7. Double-declining-balance method
    2016 $
    2017 $
  8. Activity method
    2016 $
    2017 $
Answer #1

a)

Straight line Method
Cost of Truck $     35,400
Less: Salvage value $     (4,560)
Depreciable value $     30,840
Depreciation per year ($30,840/5) $        6,168
Number of months used in 2016 9 Months
Depreciation for 2016 ($6,168/12*9) $        4,626
Depreciation for 2017 $        6,168

b)

Activity Based Method (Units of output)
Cost of Truck $ 35,400
Less: Salvage value $ (4,560)
Depreciable value $ 30,840
Total expected miles    157,200
Depreciation per mile (30,840/157,200) $      0.20
Depreciation for 2016 (9,000*$0.20) $    1,800
Depreciation for 2017 (10,300*$0.20) $    2,060

c)

Depreciation under DDB Method
Year - a Net Book value, beginning of year - b Double declained depreciation - c = b/Life of assets*2 Net book value, End of the year - d = b-c
2016 $ 35,400 $35,400/5*2*9/12 = $10,620 $      24,780
2017 $ 24,780 $                                             9,912 $      14,868

d)

Year Cost of the assets Salvage value Depreciable value sum of the year 2016 Depreciation 2017 Depreciation
2016 $                   35,400 $              4,560 $                     30,840 1+2+3+4+5 = 15 $30,840*(5/15) * (9 /12) = $7,710 $30,840*(5/15) * (3 /12) = $2,570
2017 $                   35,400 $              4,560 $                     30,840 1+2+3+4+5 = 15 $30,840*(4/15) * (9 /12) = $6,168
Depreciation expenses 7710 8738

e)

Year Book Value
2016 ($35,400-$4,626) $      30,774
2017 ($30,774-$6,168) $      24,606

f)

Year Book Value
2016 ($35,400-$1,800) $      33,600
2017 ($33,600-2,060) $      31,540

g)

Net book value, End of the year - d = b-c
$                           24,780
$                           14,868

h)

Year Depreciation expense Book Value
2016 7710 27690
2017 8738 18952

You can reach me over comment box if you have any doubts. Please rate this answer